OpenStudy (anonymous):

Assume batteries and generators provide the stated potential difference for any current that flows. A 10-Ω resistor, a 15-Ω resistor, and a 5-Ω resistor are connected in a series across a 90-V battery. a. What is the equivalent resistance of the circuit? b. What is the current in the circuit?

OpenStudy (anonymous):

So it's |---( 10-Ω )----| | ( 15-Ω ) (90V) | | | |-----( 5-Ω )---| The resistance in the circuit would just be 10-Ω + 15-Ω + 5-Ω = 30-Ω Because their in series. For the other question: How to calculate voltage drop across series resistors. Determine current of total resistors across the series. (I = Current/V=Voltage/R(t)=Resistance Total) I = V / R(t) I = 90V / (30-Ω) I = 3 A Now that we have the Current across the resistors we can calculate the voltage across EACH resistor. I = Current Vx = Voltage (x = Current Value across X Resistor) Rx = Resistor value(x = resistor#) V = I x Rx V1 = 3 A x 10 V1 = 30V V2 = 3 A x 15 V2 = 45V V3 = 3 A x 5 V3 = 15V V1 + V2 + V3 = 30V + 45V + 15V = 90V =Vtotal